OpenStudy (ray_masters):

Need help ASAP! will medal and fan! I’m not that good in math and I have no idea how to solve this :’( 5x + 3(x – 5) = 6x + 2x – 15

OpenStudy (jakesaurus1127):

You need to use order of operations. Do you know PEMDAS

OpenStudy (ray_masters):

no i don't think so i might

OpenStudy (fairegaming):

5x + 3x - 15 = 6x + 2x - 15 pair them up on each side now.

OpenStudy (jakesaurus1127):

Parentheses Exponents Multiply Divide Add Subtract

OpenStudy (fairegaming):

Jake that's wrong for this type of problem

OpenStudy (jakesaurus1127):

I was just thnking start with the parentheses

OpenStudy (ray_masters):

do we times that by 5x +3? i'm confused

OpenStudy (fairegaming):

With 3(x-5) yes you multiply but no need for pemdas

OpenStudy (ray_masters):

k so times those together, and then what?

OpenStudy (fairegaming):

Jake I already told you what to do... Take off the problem I finished for you, and pair them up ex. 5x + 3x

OpenStudy (fairegaming):

8x - 15 = 8x - 15 so it's a pair, so I am thinking the problem is correct/equal.

OpenStudy (fairegaming):

Your welcome...

OpenStudy (ray_masters):

yea thanks 4 ur help

OpenStudy (fairegaming):

I got nothing better on a school bus to do XD

OpenStudy (ray_masters):

lol

OpenStudy (fairegaming):

Your welcome...

OpenStudy (fairegaming):

8x - 15 = 8x - 15 so it's a pair, so I am thinking the problem is correct/equal.